OpenStudy (anonymous):

I need to find the domain and range, the x and y intercepts and the horizontal and vertical asymptoteand then I need to graph it. x^2+x-2/x^2-3x-4

OpenStudy (anonymous):

domain: set the denominator equal to zero and solve then say "all real numbers except those"

OpenStudy (anonymous):

shouldn't be too bad for this one since the denominator factors \[x^2-3x-4=0\] \[(x-4)(x+1)=0\] \[x=4, x=-1\] so all numbers except those

OpenStudy (anonymous):

Okay:) that seems simple enough.

OpenStudy (anonymous):

that also answers the "vertical asymptote" question vertical asymptotes are \(x=-1\) and \(x=4\)

OpenStudy (anonymous):

kill two birds with on stone on that one

OpenStudy (anonymous):

Great!

OpenStudy (anonymous):

horizontal asymptote is easy as well the degree of the numerator is 2 the degree of the denominator is 2 since the degrees are the same, the horizontal asymptote is \(y=\text{ratio of leading coefficients}\)

OpenStudy (anonymous):

in your example it is \(y=1\)

OpenStudy (anonymous):

\(x\) intercepts, set \[x^2+x-2=0\] and solve this one factors too \[(x-1)(x+2)\]

OpenStudy (anonymous):

so \(x\) intercepts are \((1,0)\) and \((-2,0)\)

OpenStudy (anonymous):

now to graph by hand, make your horizontal asymptotes and your vertical asymptote and then plot some points, not too many

OpenStudy (anonymous):

I thought that was the horizontal asymptote

OpenStudy (anonymous):

\(y=1\) is the horizontal asymptote it is a horizontal line 

OpenStudy (anonymous):